# Differentially Private Stream Processing at Scale

### v2

@@ Abstract
**footnotetext: Contributed equally.
1 Introduction
Analysis of streaming data with differential privacy (DP) [16] has been studied from the initial days of the field [17, 11], and this has been followed up in a sequence of works that include computing simple statistics [38], to machine learning applications (a.k.a. online learning) [28, 41, 2, 29, 50]. While all of these works focus on the abstract algorithmic design for various artifacts of streaming data processing, to the best of our knowledge, none of them focus on designing a scalable stream processing system. In this work, we primarily focus on designing a scalable DP stream processing system, called Differential Privacy SQL Pipelines (DP-SQLP), and make algorithmic advances along the way to cater to the scalability needs of it. DP-SQLP is implemented using a streaming framework similar to Spark streaming [49], and is built on top of the Spanner database [13] and F1 query engine [40] from Google. We also present production applications with two use cases in Section 6. The first is a real world use case that deploys DP-SQLP in Google Shopping to generate streaming page-view counts. The second applies the streaming DP algorithm to Google Trends.
@@ 1 Introduction
  4. 4.
    Effective user contribution bounding: In real applications, each person may contribute multiple records to the data stream at different times. Providing “event-level DP” (where a single action by a person/user is protected) does not provide sufficient privacy protection. In this work, we provide “user-level DP”, where all the actions by a person/user is protected simultaneously. To provide user-level DP, one has to bound the contribution of each user, and that eventually introduces bias on the statistics that get computed. But contribution bounding controls the variance of the noise we must add to ensure DP. A natural (and unavoidable [5, 36, 31]) challenge is to decide on the level of contribution bounding to balance this bias and variance.
  5. 5.
@@
Motivation for DP-SQLP: Data streams appear commonly in settings like Web logs, spatio-temporal GPS traces [6], mobile App usages [33], data generated by sensor networks and smart devices [35], live traffic in maps [47], cardiovascular monitoring [44], real-time content recommendation [39], and pandemic contact tracing [12]. Almost all of these applications touch sensitive user data on a continual basis. Calandrino et al. [8] demonstrated that continuous statistic release about individuals can act as a strong attack vector for detecting the presence/absence of a single user (in the context of collaborative recommendation systems). Hence, it is imperative to a streaming system to have rigorous privacy protections. In this work we adhere to differential privacy. For more discussion on the type of streams we consider, see a detailed survey in [27].
@@
  • Private key selection: A priori, the set of possible keys is unbounded, so our system must identify a set of relevant keys to track. To protect privacy, we cannot identify a particular key  based on the contributions of a single user. The streaming setting adds two additional complications: (a) The existence of each key  is only known when it is observed in a data record, and (b) the privacy leakage due to continually releasing information about any particular key increases the DP cost due to composition [19]. To address these challenges, we design a novel algorithm (Algorithm 1) that couples “binary tree aggregation” [17, 11, 26] (a standard tool for continual release of DP statistics that only accumulates privacy cost that is poly-logarithmic in the number of aggregate releases from the data stream) with a thresholding scheme that allows one to only operate on keys that appear in at least $\mu>0$ user records. To further minimize privacy leakage, we employ a variance reduced implementation of the binary tree aggregation protocol [26].
@@ 1.1 Problem Statement
When querying a growing database or data stream, the above-mentioned query only shows a snapshot at certain date. In stream data processing, we use event-time window to determine which records in the  event time domain to process, and triggers to define when in the  processing time domain the results of groupings are emitted [3]. Let $W$ denote the event-time windows, $W=[w_{1},w_{2}...)$. Each event-time window is defined by a starting time and an end time $w_{i}=(t_{i,s},t_{i,e})$. $D_{w_{i}}\in D$ contains all records that can be assigned to window $w_{i}$ so that the timestamp $t$ of each record satisfies $t_{i,s}\leq t<t_{i,e}$. Let $Tr$ denote a set of triggering times in the processing domain, $\texttt{Tr}=[tr_{1},tr_{2}...)$. We assume triggering time is predefined and independent to dataset [^3] In practical application, the streaming system may choose trigger adaptively, with complicated implementation Akidau et al. [3]. The streaming system incrementally processes $D_{w_{i}}$ at triggering time $_{w_{i}}=\{tr_{i,s},...,tr_{i,e}\}\subset\texttt{Tr}$ Tr. Due to the time domain skew[3], $t_{i,s}\leq tr_{i,s}$ and $t_{i,e}\leq tr_{i,e}\leq t_{i,e}+t$, where $t$ is the maximum delay that system allows for late arriving records. Our goal is to release the histogram for all sub-stream $D_{w_{i}}$, at every triggering time $tr_{i}\in\texttt{Tr}_{w_{i}}$, in a differentially private (DP) manner (See Appendix B for a formal DP definition). For a pictorial representations of various timing concepts, see Figure 1.
Figure 1: Event time domain and processing time domain
@@
Privacy implication of input driven stream: In terms of privacy, we want to ensure that the stream processing system ensures $(\varepsilon,\delta)$-user level DP [16, 15, 19] over the complete stream. Since we are operating under the constraint that the data stream is an input driven stream (Definition A.1), the timings of the system (e.g., event time (Definition A.2), processing time (Definition A.3), and triggering time (Definition A.5)) can only be defined w.r.t. times at which the inputs have appeared. Thus it forces us to define the DP semantics which considers the triggering times to be fixed across neighboring data sets (in the context of traditional DP semantics). For a given user, what we protect via DP is the actual data that is contributed to the data stream. We provide a formalism in Appendix B.
1.2 Related Work
@@
Stream processing has been an active research field for more than 20 years [24]. It is now considered to be a mature technology with various streaming frameworks deployed at scale in industry, including Spark Streaming [49], Apache Beam [3] and Apache Flink [9]. However, none of these systems offer differentially private streaming queries.
@@
Our work builds on a long line of DP research that focuses on extending one shot applications of DP mechanisms to the continual observation (streaming) setting for both analytics and learning applications [17, 11, 26, 10, 29, 34, 14, 25, 22]. These mechanisms crucially leverage the tree aggregation protocol [17, 11, 26] or variants of it based on the matrix factorization mechanism [34, 14, 25]. All these approaches have the advantage of drastically reducing the error induced by repeated application of a DP mechanism, making the DP protocol itself stateful.
@@
Of the above cited, our work is most related to [10], which investigates the problem of computing DP histograms with unknown-domains under continual observations. They leverage (extensions of) the tree aggregation protocol to build an efficient DP protocol for the continual observation setting. Our key selection algorithm (Algorithm 1) is heavily inspired by [10], with the main difference being the use of additional threshold $\mu$ for further privacy protection, and an algorithm to predict whether to select a key even if it has zero records (Algorithm 3). Algorithm 3 is crucial to the scalability of our approach to production workloads. From a system design point, our work further extend [10] as follows:
  1. 1.
@@
Another prior system that continuously releases streaming user count at scale is Google FLEDGE k-Anonymity server [21, 22]. It determines whether a given advertisement has been shown to at least k users over a sliding window with event-level differential privacy guarantee. While our algorithms and systems are more general that can be applied to arbitrary histogram query at scale, with a user-level guarantee.
1.3 Organization
@@
2 Preliminaries
In this section, we describe the formalism that will be necessary for the rest of the paper: a) DP on Streams (Appendix B), b) DP continual observation (Appendix C), and c) System architecture for the streaming system (Section 2.1). For the purpose of brevity, we will defer (a) and (b) to the appendix.
2.1 Streaming System Architecture
Figure 2: High-level overview for the DP-SQLP system
The streaming differential privacy mechanism described in this paper can be generally applied to various streaming frameworks, including Spark Streaming[49], Apache Beam[3] and Apache Flink[9]. The DP-SQLP system we develop is implemented using a streaming framework similar to Spark Streaming[49], as shown in Figure 2.
@@ 2.1 Streaming System Architecture
When processing a micro-batch, the DP-SQLP operator will interact with the system state store for a state update. Once the differentially private histogram is generated, it will be materialized to the data sink [^4] Data sink is the storage system used to store and serve output data, including file systems and databases.. Similar to Spark Streaming, our streaming framework provides consistent, ”exactly-once” processing semantic across multiple data centers. In addition, the streaming framework also provides fault tolerance and recovery.
There are multiple ways to schedule micro-batches based on certain rules[3], like processing timer based trigger, data arrival based trigger and combinations of multiple rules.
@@
Figure 3: Two types of streaming scheduling
@@
As Akidau pointed out in the unified dataflow model [3], batch, micro-batch, and pure streaming are implementation details of the underlying execution engine. Although the streaming differential privacy mechanism discussed in this paper is executed by a streaming system based on micro-batch, the mechanism and algorithm can be widely applied to batch, micro-batch, and pure streaming systems.
@@
3 Streaming Private Mechanism
Figure 4: Overview of streaming differential privacy mechanism
@@ 3 Streaming Private Mechanism
3.1 Non-Private Streaming Aggregation
The traditional streaming aggregation operator without differential privacy is shown on the top of Figure 4. Records within the micro-batch are grouped by key and aggregated by the reduce function [49]. After that, the partial aggregation result will be merged with the previous state [45] and the update histogram is emitted. There is no differential privacy protection within this process, and user privacy can be leaked from multiple dimensions, including histogram value, aggregation key and the differences between two histogram updates.
3.2 User Contribution Bounding
@@
Some one-shot mechanisms perform user contribution bounding by limiting contributed value per key and the number of contributed keys per user [4, 48]. However, this approach does not fit the streaming setting, since it requires three shuffle stages - shuffle by user, shuffle by (key, user) and shuffle by key.
User contribution bounding in streaming DP is performed on the user level for the entire data stream $D$ [^5] In a production streaming system, the DP guarantee is commonly defined with the minimum privacy protection unit (e.g., [user, day]). The maximum number of record per user $C$ needs to be enforced within each privacy unit.:
@@
Choosing the right contribution bounding parameters $C$ and $L_{m}$ is critical for privacy-utility trade-off. When the bounding limit is small, the noise is small, but the data loss may be significant (e.g. if most/all users have a lot of data to contribute). On the other side, when the bounding limit is large, the data loss is small, but noise is large. It is possible to find a near optimal point from a heuristic study, which we discuss in Section 5. Indeed, one approach to choosing a good a contribution bound is to inspect the data distribution. For example, we can pick $C$ at $99^{th}$ percentile of per-user records (i.e. $<1\%$ users have more than C records), which can be chosen in a DP way if computed on a fraction of the data stream, or in a non-DP way if it is based on proxy data.
@@
3.3 Streaming Private Key Selection
The main objective we consider in this section is to select the set of keys that exceed a certain threshold of user contributions $\mu\geq 0$. (These are the keys that are used later for releasing the aggregation columns in Section 3.4 below.) Recall that our streaming system is input driven by a growing data stream, meaning, one only sees the existence of a key if it has at least one user contribution. This poses a significant privacy challenge since the detectable set of keys is highly dependent on the data set. As a result, we design a novel thresholding scheme coupled with binary tree aggregation [17, 11, 26] that allows one to only operate with keys that deterministically have at least $\mu$ user contributions, and still preserve $(\varepsilon,\delta)$-DP. In the following, we provide a description of the algorithm, along with the privacy analysis.
Data preprocessing: After user contribution bounding (discussed in Section 3.2), we first perform a regular key GROUP BY and aggregation for all records within the current micro-batch [^6] This is a system level operation without any implication to the privacy guarantee.. Then we merge the aggregated data on each key with a data buffer that is stored in a system state [^7] For every key, we accumulate aggregation column values, as well as the number of unique users. This process is called data accumulation.. Beyond that we execute the key selection algorithm described in Algorithm 1.
@@ 3.3 Streaming Private Key Selection
Our approach is an extension of thresholding algorithm in [32] to the streaming setting. In short, we carefully select a threshold and compute (with DP noise) the number of unique users that contribute to each encountered key. If this noisy number is greater than or equal to the chosen threshold, the key is released. We describe the algorithm in full detail in Algorithm 1, and provide the formal privacy guarantee in Theorem 3.1.
@@
Remark: For brevity, in Theorems 3.1 and 3.2, we provide the guarantees assuming each user only contributes once, i.e., in the language of Section 3.2 we assume that that user contribution bound $C=1$. However, in our implementation we do allow $C>1$. The idea is to use a tighter variant of advanced composition for $(\varepsilon,\delta)$-DP [19] while ensuring that each user contributes at most once to each key in any instance of Algorithm 1. In the following we provide the privacy and the utility guarantees. For brevity, we defer the proofs to Appendix E and Appendix F.

### v3

@@ Abstract.
$\dagger$\$$\dagger$\$footnotetext: Contributed equally.
PVLDB Reference Format:
Bing Zhang, Vadym Doroshenko, Peter Kairouz, Thomas Steinke, Abhradeep Thakurta, Ziyin Ma, Eidan Cohen, Himani Apte, Jodi Spacek. PVLDB, 17(12): XXX-XXX, 2024.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 17, No. 12 ISSN 2150-8097.
doi:XX.XX/XXX.XX
1. Introduction
Analysis of streaming data with differential privacy (DP) (Dwork et al., 2006b) has been studied from the initial days of the field (Dwork et al., 2010a; Chan et al., 2011), and this has been followed up in a sequence of works that include computing simple statistics (Perrier et al., 2018), to machine learning applications (a.k.a. online learning) (Jain et al., 2012; Smith and Thakurta, 2013; Agarwal and Singh, 2017; Kairouz et al., 2021; Úlfar Erlingsson et al., 2019). While all of these works focus on the abstract algorithmic design for various artifacts of streaming data processing, to the best of our knowledge, none of them focus on designing a scalable stream processing system. In this work, we primarily focus on designing a scalable DP stream processing system, called Differential Privacy SQL Pipelines (DP-SQLP), and make algorithmic advances along the way to cater to the scalability needs of it. DP-SQLP is implemented using a streaming framework similar to Spark streaming (Zaharia et al., 2013), and is built on top of the Spanner database (Corbett et al., 2013) and F1 query engine (Samwel et al., 2018) from Google. We also present production applications with two use cases in Section 6. The first is a real world use case that deploys DP-SQLP in Google Shopping to generate streaming page-view counts. The second applies the streaming DP algorithm to Google Trends.
@@ 1. Introduction
  4. (4)
    Effective user contribution bounding: In real applications, each person may contribute multiple records to the data stream at different times. Providing “event-level DP” (where a single action by a person/user is protected) does not provide sufficient privacy protection. In this work, we provide “user-level DP”, where all the actions by a person/user is protected simultaneously. To provide user-level DP, one has to bound the contribution of each user, and that eventually introduces bias on the statistics that get computed. But contribution bounding controls the variance of the noise we must add to ensure DP. A natural (and unavoidable (Amin et al., 2019; Liu et al., 2022; Kamath et al., 2023)) challenge is to decide on the level of contribution bounding to balance this bias and variance.
  5. (5)
@@
Motivation for DP-SQLP: Data streams appear commonly in settings like Web logs, spatio-temporal GPS traces (Andrés et al., 2013), mobile App usages (Latif et al., 2020), data generated by sensor networks and smart devices (Liu et al., 2018), live traffic in maps (Wang et al., 2016), cardiovascular monitoring (Tan et al., 2021), real-time content recommendation (Phelan et al., 2009), and pandemic contact tracing (Cho et al., 2020). Almost all of these applications touch sensitive user data on a continual basis. Calandrino et al. (2011) demonstrated that continuous statistic release about individuals can act as a strong attack vector for detecting the presence/absence of a single user (in the context of collaborative recommendation systems). Hence, it is imperative to a streaming system to have rigorous privacy protections. In this work we adhere to differential privacy. For more discussion on the type of streams we consider, see a detailed survey in (Isah et al., 2019).
@@
  • Private key selection: A priori, the set of possible keys is unbounded, so our system must identify a set of relevant keys to track. To protect privacy, we cannot identify a particular key  based on the contributions of a single user. The streaming setting adds two additional complications: (a) The existence of each key  is only known when it is observed in a data record, and (b) the privacy leakage due to continually releasing information about any particular key increases the DP cost due to composition (Dwork and Roth, 2014). To address these challenges, we design a novel algorithm (Algorithm 1) that couples “binary tree aggregation” (Dwork et al., 2010a; Chan et al., 2011; Honaker, 2015) (a standard tool for continual release of DP statistics that only accumulates privacy cost that is poly-logarithmic in the number of aggregate releases from the data stream) with a thresholding scheme that allows one to only operate on keys that appear in at least $\mu>0$ user records. To further minimize privacy leakage, we employ a variance reduced implementation of the binary tree aggregation protocol (Honaker, 2015).
@@ 1.1. Problem Statement
Figure 1. Event time domain and processing time domain
[…]
@@
Privacy implication of input driven stream: In terms of privacy, we want to ensure that the stream processing system ensures $(\varepsilon,\delta)$-user level DP (Dwork et al., 2006b; Dwork et al., 2006a; Dwork and Roth, 2014) over the complete stream. Since we are operating under the constraint that the data stream is an input driven stream (Definition A.1), the timings of the system (e.g., event time (Definition A.2), processing time (Definition A.3), and triggering time (Definition A.5)) can only be defined w.r.t. times at which the inputs have appeared. Thus it forces us to define the DP semantics which considers the triggering times to be fixed across neighboring data sets (in the context of traditional DP semantics). For a given user, what we protect via DP is the actual data that is contributed to the data stream. We provide a formalism in Appendix B.
Figure 2. High-level overview for the DP-SQLP system
1.2. Related Work
@@
Stream processing has been an active research field for more than 20 years (Fragkoulis et al., 2020). It is now considered to be a mature technology with various streaming frameworks deployed at scale in industry, including Spark Streaming (Zaharia et al., 2013), Apache Beam (Akidau et al., 2015) and Apache Flink (Carbone et al., 2015). However, none of these systems offer differentially private streaming queries.
@@
Our work builds on a long line of DP research that focuses on extending one shot applications of DP mechanisms to the continual observation (streaming) setting for both analytics and learning applications (Dwork et al., 2010a; Chan et al., 2011; Honaker, 2015; Cardoso and Rogers, 2022; Kairouz et al., 2021; Li et al., 2015; Denisov et al., 2022; Henzinger and Upadhyay, 2022; Epasto et al., 2023b). These mechanisms crucially leverage the tree aggregation protocol (Dwork et al., 2010a; Chan et al., 2011; Honaker, 2015) or variants of it based on the matrix factorization mechanism (Li et al., 2015; Denisov et al., 2022; Henzinger and Upadhyay, 2022). All these approaches have the advantage of drastically reducing the error induced by repeated application of a DP mechanism, making the DP protocol itself stateful.
@@
Of the above cited, our work is most related to (Cardoso and Rogers, 2022), which investigates the problem of computing DP histograms with unknown-domains under continual observations. They leverage (extensions of) the tree aggregation protocol to build an efficient DP protocol for the continual observation setting. Our key selection algorithm (Algorithm 1) is heavily inspired by (Cardoso and Rogers, 2022), with the main difference being the use of additional threshold $\mu$ for further privacy protection, and an algorithm to predict whether to select a key even if it has zero records (Algorithm 3). Algorithm 3 is crucial to the scalability of our approach to production workloads. From a system design point, our work further extend (Cardoso and Rogers, 2022) as follows:
  1. (1)
@@
Another prior system that continuously releases streaming user count at scale is Google FLEDGE k-Anonymity server (Epasto et al., 2023a, b). It determines whether a given advertisement has been shown to at least k users over a sliding window with event-level differential privacy guarantee. While our algorithms and systems are more general that can be applied to arbitrary histogram query at scale, with a user-level guarantee. An extended list of real-world uses of DP is presented in Desfontaines (2024).
1.3. Organization
@@
2. Preliminaries
In this section, we describe the formalism that will be necessary for the rest of the paper: a) DP on Streams (Appendix B), b) DP continual observation (Appendix C), and c) System architecture for the streaming system (Section 2.1). For the purpose of brevity, we will defer (a) and (b) to the appendix. For a comprehensive introduction to DP, please refer to (Vadhan, 2017).
2.1. Streaming System Architecture
The streaming differential privacy mechanism described in this paper can be generally applied to various streaming frameworks, including Spark Streaming(Zaharia et al., 2013), Apache Beam(Akidau et al., 2015) and Apache Flink(Carbone et al., 2015). The DP-SQLP system we develop is implemented using a streaming framework similar to Spark Streaming(Zaharia et al., 2013), as shown in Figure 2.
@@ 2.1. Streaming System Architecture
[…]
There are multiple ways to schedule micro-batches based on certain rules(Akidau et al., 2015), like processing timer based trigger, data arrival based trigger and combinations of multiple rules.
@@
Figure 3. Two types of streaming scheduling
@@
As Akidau pointed out in the unified dataflow model (Akidau et al., 2015), batch, micro-batch, and pure streaming are implementation details of the underlying execution engine. Although the streaming differential privacy mechanism discussed in this paper is executed by a streaming system based on micro-batch, the mechanism and algorithm can be widely applied to batch, micro-batch, and pure streaming systems.
@@
Figure 4. Overview of streaming differential privacy mechanism
3. Streaming Private Mechanism
@@
3.1. Non-Private Streaming Aggregation
The traditional streaming aggregation operator without differential privacy is shown on the top of Figure 4. Records within the micro-batch are grouped by key and aggregated by the reduce function (Zaharia et al., 2013). After that, the partial aggregation result will be merged with the previous state (To et al., 2018) and the update histogram is emitted. There is no differential privacy protection within this process, and user privacy can be leaked from multiple dimensions, including histogram value, aggregation key and the differences between two histogram updates.
3.2. User Contribution Bounding
@@
Some one-shot mechanisms perform user contribution bounding by limiting contributed value per key and the number of contributed keys per user (Amin et al., 2022; Wilson et al., 2019). However, this approach does not fit the streaming setting, since it requires three shuffle stages - shuffle by user, shuffle by (key, user) and shuffle by key.
[…]
@@
Choosing the right contribution bounding parameters $C$ and $L_{m}$ is critical for privacy-utility trade-off. When the bounding limit is small, the noise is small, but the data loss may be significant (e.g. if most/all users have a lot of data to contribute). On the other side, when the bounding limit is large, the data loss is small, but noise is large. It is possible to find a near optimal point from a heuristic study, which we discuss in Section 5. Indeed, one approach to choosing a good contribution bound is to inspect the data distribution. For example, we can pick $C$ at $99^{th}$ percentile of per-user records (i.e. $<1\%$ users have more than C records), which can be chosen in a DP way if computed on a fraction of the data stream, or in a non-DP way if it is based on proxy data.
@@
3.3. Streaming Private Key Selection
The main objective we consider in this section is to select the set of keys that exceed a certain threshold of user contributions $\mu\geq 0$. (These are the keys that are used later for releasing the aggregation columns in Section 3.4 below.) Recall that our streaming system is input driven by a growing data stream, meaning, one only sees the existence of a key if it has at least one user contribution. This poses a significant privacy challenge since the detectable set of keys is highly dependent on the data set. As a result, we design a novel thresholding scheme coupled with binary tree aggregation (Dwork et al., 2010a; Chan et al., 2011; Honaker, 2015) that allows one to only operate with keys that deterministically have at least $\mu$ user contributions, and still preserve $(\varepsilon,\delta)$-DP. In the following, we provide a description of the algorithm, along with the privacy analysis.
[…]
@@ 3.3. Streaming Private Key Selection
Our approach is an extension of thresholding algorithm in (Korolova et al., 2009) to the streaming setting. In short, we carefully select a threshold and compute (with DP noise) the number of unique users that contribute to each encountered key. If this noisy number is greater than or equal to the chosen threshold, the key is released. We describe the algorithm in full detail in Algorithm 1, and provide the formal privacy guarantee in Theorem 3.1.
@@
Remark: For brevity, in Theorems 3.1 and 3.2, we provide the guarantees assuming each user only contributes once, i.e., in the language of Section 3.2 we assume that that user contribution bound $C=1$. However, in our implementation we do allow $C>1$. The idea is to use a tighter variant of advanced composition for $(\varepsilon,\delta)$-DP (Dwork and Roth, 2014) while ensuring that each user contributes at most once to each key in any instance of Algorithm 1. In the following we provide the privacy and the utility guarantees. For brevity, we defer the proofs to Appendix E and Appendix F.